OpenStudy (anonymous):

Simplify: 24/4*3-6*5

OpenStudy (anonymous):

I got -4/3

OpenStudy (anonymous):

what I did is first I multiplied 4 by 3=12 and 6 by 5= 30. 12-30= -18. then 24/-18

OpenStudy (anonymous):

Apparently the answer is -12 but I don't know how that's the answer. I got -28.

OpenStudy (anonymous):

They must have divided 24 by 4 first in that answer.

OpenStudy (mathstudent55):

24/4*3-6*5 First, you need to do multiplications and divisions in the order they appear from left to right. = 6 * 3 - 30 = 18 - 30 Now do the subtraction. = -12

OpenStudy (anonymous):

Isnt there a rule where you are supposed to work left to right with this kind of order of operations?

OpenStudy (anonymous):

nice.

OpenStudy (anonymous):

PEMDAS?

OpenStudy (mathstudent55):

Yes, that's what I did, but it's not simply left to right. If you only have multiplications, divisions, additions, and subtractions, first you do the multiplications and divisions in the order they appear from left to right. After all multiplications and divisions are done, then you move on to all additions and subtractions in the order they appear from left to right.

OpenStudy (mathstudent55):

Yes, PEMDAS, but I am only using the MDAS part here.

OpenStudy (mathstudent55):

PEMDAS itself needs an explanation. P - Parentheses come first E - exponents next MD - means multiplications & divisions in the order they appear from left to right AS - additions & subtractions are also from left to right in the order they appear.

OpenStudy (anonymous):

ah k

OpenStudy (anonymous):

I need help with one similar